OpenStudy (anonymous):

write an equation(slope): m=8/9,(6,-3)

OpenStudy (aroub):

y=mx+b , m=8/9 , (6,-3) -3=8/9(6)+b -3=16/3+b -3-16/3=b -9/3-16/3=b -25/3=b so the equation is, y=8/9x-25/3